In a series of 2ohms,4ohms,3ohms,2ohms,4ohms and 3voltage find the current in the circuit


1
Expert's answer
2022-02-16T08:35:00-0500

The Ohm's law says

"I=\frac{V}{R_{\rm net}}"

In our case

"I=\frac{3\:\rm V}{\rm(2+4+3+2+4)\:\Omega}=0.2\:\rm A"